# https://re-ground.github.io

### v2

@@ 6 Experiments
Figure 5: Qualitative comparisons. Stable Diffusion (SD, 2nd column) generates images that align with the given text descriptions, including the underlined phrase in each row, but cannot take the bounding box constraints as input. GLIGEN (3rd column) creates images that match the input layouts but suffers from the description omission issue, failing to include the underlined descriptions in the final image. The scheduled sampling strategy (4th column) can partially address this issue (for instance, in the 3rd row, where “airline status board” appears behind the clock), but it results in a noticeable decline in spatial accuracy (as seen in the 1st row, where the tie is not generated). In contrast, our method (last column) accurately incorporates the underlined text descriptions while maintaining precise spatial representation.
@@ 6.1 Datasets
The validation set of the MS-COCO-2017 dataset [28] consists of 5,000 image-annotation pairs, each of which includes the $xy$-coordinate of each bounding box and the corresponding object category. As GLIGEN [26] is trained to handle a maximum of 30 bounding boxes per image, we exclude pairs that include more than 30 bounding boxes and with no bounding boxes, resulting in a total of 4,952 images. We also use the the validation set of the MS-COCO-2014 dataset [28], in which we randomly sample 5,000 pairs for evaluation.
@@
The Numerical and Spatial Reasoning (NSR-1K) benchmark, introduced by Feng et al. [10], is a collection of layout-caption pairs designed to assess the numerical and spatial reasoning capabilities of image generation methodologies. The object labels and bounding boxes are from COCO [28], while the captions are newly annotated based on the spatial relationships and numerical properties of objects. NSR-1K consists of two subsets: Counting and Spatial. We randomly sample 1,000 pairs from the Counting set and use all 1,021 pairs from the Spatial set. Then based upon the subsets, we develop a new benchmark, NSR-1K-GPT, augmenting each original caption in NSR-1K by employing ChatGPT [31] to (i) elaborate on the descriptions of each mentioned entity and (ii) provide additional details about the background of the image.
@@ 6.2 Evaluation Metrics
  • YOLO score: The spatial grounding accuracy is assessed using the YOLO score [49]. We use the pretrained YOLOv7 [49] to detect objects in each generated image and compute the average precision (AP) based on the ground truth bounding box annotations in the MS COCO dataset.
  • CLIP score: The textual grounding accuracy is assessed using CLIP score [13]. The CLIP score is measured by computing the cosine similarity between each generated image and the input text prompt.
  • FID: The image quality and diversity are evaluated using FID [14].
  • PickScore: The human preference for different methods is evaluated by PickScore [24], a CLIP-based [35] scoring function that reflects human preferences.
@@ 6.3 Comparison with GLIGEN
Figure 6: Comparison on the MS-COCO [28] dataset. Each plot shows the relationship between textual grounding (i.e. CLIP score [13]) and spatial grounding (i.e. YOLO score [49]) accuracy of GLIGEN [26] and our method. Note that the plot representing our ReGround is positioned in the top-right quadrant relative to GLIGEN, signifying that it alleviates the inherent trade-off between textual and spatial grounding.
@@
We first examine the trade-off between textual and spatial grounding for both GLIGEN [26] and our ReGround, the rewired version of GLIGEN, while varying the scheduled sampling parameter $\gamma$ from 1.0 to 0.1.
Fig. 6 presents the graphs for CLIP score [13] and YOLO score [49] measured using the MS-COCO-2014 and MS-COCO-2017 datasets. Using the MS-COCO-2014 dataset, when reducing $\gamma$ from 1.0 to 0.1, the CLIP score of GLIGEN varies from 30.44 to 31.65, while the YOLO score significantly drops from 58.13 to 22.75 (red curve in Fig. 6). In contrast, our ReGround, (blue in Fig. 6), demonstrates a notably superior trade-off between textual and spatial grounding. Specifically, with $\gamma$ set to 1.0, ReGround already achieves a CLIP score of 31.29, accounting for 70.25% of GLIGEN’s total improvement in CLIP score when $\gamma$ is reduced from 1.0 to 0.1. Despite this significant increase in CLIP score, the YOLO score remains largely unchanged, marking 56.96 which represents only a 3.31% decrease in the range of YOLO score variation for GLIGEN when $\gamma$ is adjusted from 1.0 to 0.1. Moreover, when varing the $\gamma$, the plot for ReGround (blue) is constantly on the upper right side of GLIGEN (red), signifying a more advantageous trade-off across varying $\gamma$. The same pattern is observed in the MS-COCO-2017 dataset, where our ReGround achieves 68.33% of the increase in CLIP score of GLIGEN while only compromising YOLO score by 2.62% compared to the decrease for GLIGEN.
We further show the quantitative comparisons on the two subsets of the newly generated NSR-1K-GPT benchmark in Fig. 7. The plots reveal a consistent trend with the MS-COCO datasets. For NSR-1K-GPT-Counting (Fig. 7-(a)), by reducing $\gamma$ from 1.0 to 0.1, GLIGEN’s CLIP score is increased from 32.46 to 33.67, while the YOLO score is decreased from 65.36 to 26.38. In contrast, when $\gamma=1.0$, ReGround achieves a CLIP score of 33.20, which is equal to 61.16% of GLIGEN’s total improvement in CLIP score, while the compromise in YOLO score is equal to only 3.69% of the total decrease in the YOLO score of GLIGEN from $\gamma=1.0$ to $\gamma=0.1$. For NSR-1K-GPT-Spatial (Fig. 7-(b)), the minimum CLIP score of our ReGround (33.89 at $\gamma=1.0$) is already higher than GLIGEN’s maximum CLIP score (33.88 at $\gamma=0.1$), indicating that ReGround obtains a significant enhancement in textual grounding while preserving the spatial grounding. These results highlights that the advantage of our ReGround holds robustly for the realistic image captions provided in the MS-COCO [28], as well as for diverse text prompts generated by ChatGPT [31].
Figure 7: Comparison on the two subsets of NSR-1K-GPT. Consistent with the findings from the MS-COCO [28], ReGround shows improved performance on the textual and spatial groundings, as seen by the higher CLIP score [13] for the same range of YOLO score [49].
@@
Although ReGround outperforms GLIGEN [26] on MS-COCO datasets, the layout-caption pairs in MS-COCO are not optimally designed for evaluating the description omission behavior, since we find that most of the entities in the text prompt are covered in the bounding boxes. To further assess the extent of description omission in each method, we modify the MS-COCO-2017 dataset [28] to make the COCO-Drop dataset. In this version, the bounding boxes for 50% of the categories are randomly removed from each image, thereby preventing every entity described in the text prompt from being included within the bounding boxes.
Fig. 9 shows the quantitative evaluation results of ReGround and GLIGEN on COCO-Drop. In this case, ReGround shows a larger advantage over GLIGEN in CLIP score, obtaining a gap in CLIP score which is 1.57 times that of the original MS-COCO-2017 dataset before box dropping for $\gamma=1.0$. Such a larger gap in CLIP score demonstrates that compared to GLIGEN, our ReGround better reflects the text prompts even when some entities in the text prompt are not provided as a bounding box. Fig. 10 displays a representative example, where GLIGEN fails to generate a cupcake when its corresponding bounding box is removed in COCO-Drop, whereas our ReGround robustly generates the cupcake even when it it not provided as a bounding box.
@@
Fig. 9 displays the relationship between YOLO score [49] and FID [14] for each method on MS-COCO-2017. Note that the FID of ReGround is constantly lower than that of GLIGEN [26], meaning that our network rewiring also results in higher image quality.
@@
We measure the user preference of the images given a text prompt for our ReGround over GLIGEN [26] through PickScore [24]. On MS-COCO-2017, ReGround is preferred over GLIGEN by 55.66% to 44.34%, and on COCO-Drop, ReGround is preferred by 57.57% to 42.43%.
Figure 11: Comparison of applying BoxDiff [50] on GLIGEN [26] and our ReGround, respectively. (a) and (b) show that our ReGround further improves the grounding quality of BoxDiff on NSR-1K-GPT datasets. (c) While BoxDiff with GLIGEN (left) also shows description omission – omitting “beautiful sunset” from the text prompt – BoxDiff with our ReGround contains the sunset in the final image (right).
@@ 6.4 Impact of ReGround as a Backbone
We demonstrate that applying our rewiring of attention modules can also improve text-image alignment in other layout-guided generation methods that use GLIGEN as a backbone. For instance, BoxDiff [50] is a notable example that uses GLIGEN as its foundation and improves spatial grounding with respect to the bounding boxes by leveraging cross-attention maps as additional spatial cues in a zero-shot manner. Our network rewiring can also be combined with the zero-shot approach of BoxDiff. Fig. 11 illustrates the results on the NSR-1K-GPT datasets (a) when BoxDiff uses GLIGEN as the base, and (b) when it uses our ReGround, the rewired GLIGEN, as the base. It depicts that for the same range of spatial grounding accuracies, ReGround obtains noticeably higher textual grounding (i.e. CLIP score [13]). Also, as shown in Fig. 11-(c), our network rewiring allows for a more detailed description to accurately appear in the final image, both for the entities in the bounding boxes (“truck”) and the entities that are given as a text prompt (“sunset”).

### v3

@@ 6 Experiments
Figure 5: Qualitative comparisons. Stable Diffusion (SD, 2nd column) generates images that align with the given text descriptions, including the underlined phrase in each row, but cannot take bounding boxes as input. GLIGEN (3rd column) creates images that match the input layouts but suffers from description omission, failing to reflect the underlined descriptions. Scheduled sampling strategy (4th column) can partially address this issue (for instance, in the 5th row, where “window” appears in the room), but it results in a noticeable decline in spatial accuracy (as seen in the 1st row, where the tie is not generated). In contrast, our method (last column) accurately incorporates the underlined text descriptions while maintaining precise spatial representation.
@@ 6.1 Datasets
We use the validations sets of both MS-COCO-2014 and MS-COCO-2017 datasets [31]. Each dataset provides image-captions pairs and the $xy$-coordinates of bounding boxes along with their corresponding object categories.
@@
We also use the NSR-1K benchmark [12] for evaluation. Based on each subset of NSR-1K—Counting and Spatial—we develop a new benchmark, NSR-1K-GPT, augmenting each original caption in NSR-1K using GPT-4 [35]. The instructions for augmentation are to (i) elaborate on the descriptions of each mentioned entity and (ii) provide additional details about the background of the image. More details on the evaluation datasets are provided in the Appendix (Sec. 8.1).
@@ 6.2 Evaluation Metrics
  • YOLO score: Spatial grounding accuracy is assessed using YOLO score [53]. We employ YOLOv7 [53] to detect objects in each generated image and compute the average precision (AP) based on the ground truth bounding box annotations from MS-COCO [31].
  • CLIP score: Textual grounding accuracy is assessed using CLIP score [15].
  • FID: Image quality and diversity are evaluated using FID [16].
  • User Study and PickScore: We conduct a user study to assess human preferences for the generated images based on each input text prompt. Additionally, we use PickScore [26], a human preference predictor, to further analyze the results.
@@ 6.3 Comparison with GLIGEN
Figure 6: Comparisons on MS-COCO [31] and NSR-1K-GPT. Each plot shows the relationship between textual grounding (i.e. CLIP score [15]) and spatial grounding (i.e. YOLO score [53]) accuracy of GLIGEN [28] and our method. Note that the plot of our ReGround is positioned in the top-right quadrant relative to GLIGEN, signifying that it alleviates the inherent trade-off between textual and spatial grounding.
@@
We first examine the trade-off between textual and spatial groundings for both GLIGEN [28] and our ReGround, the rewired version of GLIGEN, while varying the scheduled sampling parameter $\gamma$ from 1.0 to 0.1.
Fig. 6-(a), (b) present the graphs of CLIP score [15] and YOLO score [53] measured on the MS-COCO datasets [31]. In MS-COCO-2014, when reducing $\gamma$ from 1.0 to 0.1, the CLIP score of GLIGEN varies from 30.44 to 31.65, while the YOLO score significantly drops from 58.13 to 22.75 (red in Fig. 6). In contrast, our ReGround, (blue in Fig. 6), demonstrates a notably superior trade-off between textual and spatial grounding. Specifically, with $\gamma$ set to 1.0, ReGround already achieves a CLIP score of 31.29, accounting for 70.25% of GLIGEN’s total improvement in CLIP score when $\gamma$ is reduced from 1.0 to 0.1. Despite this significant increase in CLIP score, the YOLO score remains largely unchanged, marking 56.96 which represents only a 3.31% decrease in the range of YOLO score variation for GLIGEN when $\gamma$ is adjusted from 1.0 to 0.1. Moreover, when varing the $\gamma$, the plot for ReGround (blue) is constantly on the upper right side of GLIGEN (red), signifying a more advantageous trade-off across varying $\gamma$. The same pattern is observed in MS-COCO-2017, where our ReGround achieves 68.33% of the increase in CLIP score of GLIGEN while only compromising YOLO score by 2.62% compared to the decrease for GLIGEN.
Fig. 6-(c) further shows a quantitative comparison on the Counting subset of the newly generated NSR-1K-GPT benchmark. The plot reveals a consistent trend with the MS-COCO datasets. By reducing $\gamma$ from 1.0 to 0.1, GLIGEN’s CLIP score is increased from 32.46 to 33.67, while the YOLO score is decreased from 65.36 to 26.38. In contrast, when $\gamma=1.0$, ReGround achieves a CLIP score of 33.20, which is equal to 61.16% of GLIGEN’s total improvement in CLIP score, while the compromise in YOLO score is equal to only 3.69% of the total decrease in the YOLO score of GLIGEN from $\gamma=1.0$ to $\gamma=0.1$. Moreover, a comparison on the Spatial subset of NSR-1K-GPT is provided in the Appendix (Sec. 8.2). These results highlight that the advantage of our ReGround holds robustly for the realistic image captions provided in the MS-COCO [31], as well as for diverse text prompts generated by GPT-4 [35].
@@
To further assess the extent of description omission in each method, we modify the MS-COCO-2017 dataset [31] to make the COCO-Drop dataset. In this version, the bounding boxes for 50% of the categories are randomly removed from each image, thereby preventing every entity described in the text prompt from being included within the bounding boxes.
Fig. 9 shows the quantitative comparison of ReGround and GLIGEN on COCO-Drop. In this case, ReGround shows a larger advantage over GLIGEN in CLIP score, obtaining a gap in CLIP score which is 1.57 times that of the original MS-COCO-2017 dataset before box dropping for $\gamma=1.0$. Such a larger gap in CLIP score demonstrates that compared to GLIGEN, our ReGround better reflects the text prompts even when some entities in the text prompt are not provided as a bounding box. Fig. 7 displays a representative example, where GLIGEN fails to generate a “cupcake” when its corresponding bounding box is removed in COCO-Drop, whereas our ReGround robustly generates the cupcake even when it it not provided as a bounding box.
@@
Fig. 9 displays the relationship between YOLO score [53] and FID [16] for each method on MS-COCO-2017. Note that the FID of ReGround is constantly lower than that of GLIGEN [28], meaning that our network rewiring also results in higher image quality and diversity.
User Study.
We conducted a user study to compare GLIGEN and our ReGround in terms of faithfulness to input text prompts. We used GPT-4 [35] to generate 100 prompts each containing two different objects, along with a bounding box for each object. Participants were given the text prompt along with two images—one from each method—and asked to choose the image that ‘‘better includes all the objects from the prompt.’’ Among the 92 out of 100 participants who passed the vigilance tests, our ReGround surpassed GLIGEN, with a preference rate of 70.05% compared to 29.95%. Further details on the user study are provided in the Appendix (Sec. 8.1).
@@
We further compare the PickScore [26] of GLIGEN [28] and our ReGround given each input text prompt. On MS-COCO-2017, ReGround is preferred over GLIGEN by 55.66% to 44.34%, and on COCO-Drop, ReGround is preferred by 57.57% to 42.43%.
Figure 10: Comparison of applying BoxDiff [55] on GLIGEN [28] and our ReGround, respectively. (a) and (b) show that our ReGround further improves the grounding quality of BoxDiff on NSR-1K-GPT datasets. (c) While BoxDiff with GLIGEN (left) also shows description omission—omitting “beautiful sunset” from the text prompt—BoxDiff with our ReGround contains the sunset in the final image (right).
@@ 6.4 Impact of ReGround as a Backbone
We demonstrate that applying our rewiring of attention modules can also improve text-image alignment in other layout-guided generation methods that use GLIGEN as a backbone. For instance, BoxDiff [55] is a notable example that uses GLIGEN as its foundation and improves spatial grounding with respect to the bounding boxes by leveraging cross-attention maps as additional spatial cues in a zero-shot manner. Our network rewiring can also be combined with the zero-shot guidance of BoxDiff. Fig. 10 illustrates the results on the NSR-1K-GPT datasets (a) when BoxDiff uses GLIGEN as the base, and (b) when it uses our ReGround, the rewired GLIGEN, as the base. It depicts that for the same range of spatial grounding accuracies, ReGround obtains noticeably higher textual grounding (i.e. CLIP score [15]). Also, as shown in Fig. 10-(c), our network rewiring allows for a more detailed description to accurately appear in the final image, both for the entities in the bounding boxes (“truck”) and the entities that are given as a text prompt (“sunset”).
@@ 7 Conclusion
Appendix.
Due to limited space, we provide the following contents in the Appendix: details on the evaluation setup (Sec. 8.1), additional quantitative (Sec. 8.2) and qualitative (Sec. 8.4) comparisons, and more results of ReGround as a backbone for other layout-guided generation methods (Sec. 8.3).
Acknowledgments
This work was supported by NRF grant (RS-2023-00209723), IITP grants (2022-0-00594, RS-2023-00227592, RS-2024-00399817), and Alchymist Project Program (RS-2024-00423625) funded by the Korean government (MSIT and MOTIE), and grants from the DRB-KAIST SketchTheFuture Research Center, NAVER-intel, Adobe Research, Hyundai NGV, KT, and Samsung Electronics.